# Entropy conservative discretization of compressible Euler equations with an arbitrary equation of state

### v1

@@ Abstract
This study proposes a novel spatial discretization procedure for the compressible Euler equations which guarantees entropy conservation at a discrete level when an arbitrary equation of state is assumed. The proposed method, based on a locally-conservative discretization, guarantees also the spatial conservation of mass, momentum, and total energy and is kinetic energy-preserving. In order to achieve the entropy-conservation property for an arbitrary non-ideal gas, a general strategy is adopted based on the manipulation of discrete balance equations through the imposition of global entropy conservation and the use of a summation by parts rule. The procedure, which is extended to an arbitrary order of accuracy, conducts to a general form of the internal-energy numerical flux which results in a nonlinear function of thermodynamic and dynamic variables and still admits the mass flux as a residual degree of freedom. The effectiveness of the novel entropy-conservative formulation is demonstrated through numerical tests making use of some of the most popular cubic equations of state.
@@ 1 Introduction
Numerical simulations are key in understanding the physical characteristics of turbulent compressible flows and are an invaluable tool for application-oriented studies [1]. However, it is well known that, even in the absence of shock waves, standard discretizations are susceptible to strong nonlinear instabilities at high values of the Reynolds number, especially when non-dissipative schemes are employed for the discretization of the convective terms [2]. This behavior can be traced back to the failure of the numerical method to reproduce at a discrete level some structural properties or invariants of the continuous system of equations. In fact, compressible Euler equations, which are a mathematical expression of the conservation of mass, momentum and total energy, have many induced symmetries which are typically not reproduced by the discrete system. The numerical community has progressively realized that the traditional focus on classical concepts of numerical analysis, such as order of accuracy and asymptotic convergence, has to be complemented by some more ‘physical’ concepts, such as the discrete reproduction of the basic structural properties of the continuous formulation on an arbitrary finite mesh. These methods are referred to as structure preserving or physically compatible methods and have been the subject of many studies in recent years [2, 3, 4, 5].
@@
Examples of structural properties for the compressible flow equations are the conservation of entropy (for smooth, nonviscous flows), the conservation of kinetic energy (for nonviscous flows in the incompressible limit), and the equilibrium of constant pressure and velocity distributions, which allows the existence of density waves moving at constant velocity and pressure. Several numerical methods have been developed, which are able to preserve these properties at discrete level. Kinetic Energy Preserving (KEP) methods are the most well known and studied among the others. These methods are designed to automatically preserve the conservative structure of the convective terms in the induced discrete equation for the evolution of kinetic energy. The study of KEP methods for compressible flows was pioneered by Feiereisen et al. [6]. Inspired by some formulations used for incompressible flows, the authors proposed for the first time the use of a skew-symmetric like splitting of the convective terms in mass and momentum equations to induce a conservative structure of the convective terms in the kinetic-energy equation. This work stimulated the attention on this important topic and several other studies have appeared over the years, among which we mention here the Finite Volume (FV) analyses of Jameson [7] and Subbareddy and Candler [8] and the works on the triple splitting by Kennedy and Gruber [9] and Pirozzoli [10]. More recently, also less conventional approaches have been explored, in which square-root variables are used [11, 12, 13, 14]. A quite complete characterization of the possible KEP formulations in a Finite-Difference (FD) framework has been derived only recently [15], and its relations with FV formulations have also been explored [16, 17]. As KEP methods involve a proper discretization of mass and momentum equations alone, they are agnostic to the gas model used in the integration of the total-energy equation. KEP methods have been applied mainly in the context of ideal gases, but applications to more general equations of state have shown improved robustness as well ([18, 19]).
@@
Discrete entropy conservation has also been explored to improve fidelity and suppress instabilities in turbulent simulations. Most of the contributions have been developed by following the elegant theory proposed by Tadmor [20, 21] in the context of FV methods. Entropy Conservative (EC) numerical fluxes have been derived since the early papers on this subject in the context of ideal gases [20, 21, 22]. Ismail and Roe [23] developed ‘affordable’ EC fluxes based on the logarithmic mean, whereas Chandrashekar [24] proposed for the first time both KEP and EC numerical fluxes. More recently, Ranocha [25, 26, 27] proposed a KEP and EC formulation which is also able to enforce the so-called Pressure Equilibrium Preserving (PEP) property ([28],[27], [18] ), i.e. the ability of the discrete scheme to reproduce the traveling density wave solutions of the compressible Euler equations obtained for initially uniform pressure and velocity distributions. Recently, De Michele and Coppola [29] proposed a family of Asymptotically Entropy Conservative (AEC) schemes which are PEP and KEP, and arbitrarily reduce the error on entropy conservation. The main advantage of these schemes is that, being based on arithmetic and harmonic means, they only use algebraic operations for the computation of the fluxes, and do not exhibit singularities, with increased efficiency at comparable performances. AEC schemes improve already existing formulations which approximately enforce the entropy-conservation property, as the KEEP or KEEP(N) schemes [30, 31].
@@
All the mentioned studies on EC formulations have been developed in the context of ideal gases, for which a simple linear relation between pressure and internal energy is assumed. However, numerical simulations of real gases, often in transcritical regimes, have grown significantly in recent years, due to a wide range of possible engineering applications [32, 33, 34, 19]. Despite this growing interest, the literature on structure-preserving methods explicitly designed for real gases is still scarce. Most of the studies have been focused on the pressure equilibrium property, as the generation of spurious pressure oscillations, especially in the case of multi-component flows, is a longstanding issue in scale-resolving simulations with real gases [35, 36, 37, 18]. Despite these efforts, no numerical method is known at present that can discretely preserve the primary invariants mass, momentum, and total energy and simultaneously enforce the pressure equilibrium property or the conservation of entropy in the case of gases with an arbitrary equation of state.
@@ 2 Problem formulation
where $\rho$ is the density, $p$ the pressure, $u_{\alpha}$ the Cartesian velocity component along the direction $x_{\alpha}$ and $E$ the total energy per unit mass, sum of internal and kinetic energies: $E=e+u_{\alpha}u_{\alpha}/2$. In this last equation, in Eqs. (1)-(3) and in what follows, the summation convention over repeated Greek indices is assumed. The system (1)-(3) is closed by the specification of an equation of state, which is assumed in the form $p=p(\rho,e)$ and is left as unspecified for the moment.
In this paper we focus on the discretization of the spatial terms of the system of equations (1)-(3) in the context of a semi-discretized procedure. As is usual in this framework, after the spatial discretization step has been performed, the resulting system of Ordinary Differential Equations (ODE) is integrated in time by using a standard solver, and we assume that the effects of time integration errors are negligible at sufficiently small time steps. Consequently, in our theoretical analysis we assume that all the manipulations involving time derivatives can be carried out at the continuous level.
Although we will repeatedly refer to numerical fluxes, which are usually specified in the context of FV methods, our framework is mainly that of a FD treatment of the spatial terms in Eqs. (1)–(3). Numerical fluxes arise in our analysis because we assume that the discrete formulation of the convective terms in Eqs. (1)-(3), which are directly discretized, is always locally conservative, i.e. it can be expressed as the sum of differences of numerical fluxes at adjacent faces along the three coordinate directions. A locally-conservative discretization is also globally conservative, i.e. the sum of all the pointwise contributions on the domain is zero (or a function of boundary conditions) by virtue of the telescoping property. These concepts are familiar in FV methods, but our numerical fluxes are always interpreted as functions of nodal values over a colocated uniform mesh, although a generalization of our findings to more general situations and to other discretization techniques can be easily carried out in most cases. Latin subscripts as $i,j$ or $k$ are used to denote the values of the discretized variable on a nodal point $x_{i}$.
@@
and subtracting Eq. (4) from Eq. (3) gives the evolution equation for internal energy
@@ 3 Discrete formulation
where $\mathcal{C}_{\rho\phi}$ is the convective term and $\mathcal{P}_{\rho\phi}$ is a pressure term. In Eqs. (1)–(6) $\phi$ assumes the values $1,u_{\beta},E,\kappa,e$ and $s$, respectively, whereas the pressure term is absent in mass (1) and entropy (6) equations. The symbols $\mathcal{C}$ and $\mathcal{P}$ will be used indifferently to denote both the continuous or discrete versions of the spatial terms at the right-hand sides of Eqs. (1)–(6). When needed, we will use the more explicit notation $\left.\mathcal{C}_{\rho\phi}\right|_{i}$ to denote a particular discretization of $\mathcal{C}_{\rho\phi}$ at node $x_{i}$. The theory will be developed for the one-dimensional version of Eq. (1)–(3), the extension to the three-dimensional case being straightforward, since the convective terms along the various directions are independent and can be treated separately. In this section and in the following Sec. 4 we will expose the theory with respect to second-order formulations, corresponding to two-point fluxes. In Sec. 6 the extension to high-order fluxes is carried on.
@@
where $h$ is the spatial step size and $\delta^{\,-}$ is the backward difference operator: $\delta^{\,-}a_{i}=a_{i}-a_{i-1}$. In what follows we will need to consider also the forward difference operator, which is denoted by $\delta^{\,+}$, with $\delta^{\,+}b_{i}=b_{i+1}-b_{i}$. The operators $\delta^{\,-}$ and $-\delta^{\,+}$ are one the adjoint of the other with respect to the standard scalar product. This means that, neglecting the effects of boundary conditions (or by assuming periodic boundary conditions), the following property is easily seen to hold
@@ 4.1 General treatment
A KEP discretization of mass and momentum equations can be implemented with a latitude on the specification of the mass flux $\mathcal{F}_{\rho}$. Once this has been done, the kinetic energy flux is fixed as in Eq. (10), and one is left with the internal-energy flux $\mathcal{F}_{\rho e}$ as additional residual degree of freedom to satisfy further structural properties of the formulation. In this section, we analyze the induced discrete equation for the specific entropy $\rho s$ to obtain a formulation that is globally conservative of entropy.
@@
where $g=e-Ts+p/\rho$ is the Gibbs free energy. This equation (also reported as Eq. (15.5.2) in [40], p. 352) reduces to the Eq. (33) reported in [15] in the case of ideal gases. Equation (12) is now used to obtain a relation between the discretized spatial terms of mass, internal energy and entropy. In fact, substituting the right-hand sides of Eq. (1), (5) and (6), as expressed in the generic form of Eq. (7), one has
@@
We now assume that mass and internal energy convective terms are discretized using a locally-conservative formulation, as expressed in Eq. (8): $\mathcal{C}_{\rho}=\delta^{\,-}\mathcal{F}_{\rho}/h$ and $\mathcal{C}_{\rho e}=\delta^{\,-}\mathcal{F}_{\rho e}/h$. Moreover, for the pressure term in the internal-energy equation $\mathcal{P}_{\rho e}=p\partial u/\partial x$ we use a standard, second-order central derivative formula, which can be expressed as
@@
[…]
@@
[…]
@@
Eq. (20) is the fundamental relation defining the class of EC schemes we analyze in this paper which still contains the mass flux $\mathcal{F}_{\rho}$ as a residual degree of freedom. In the present study we will only consider the scheme having as a mass flux simply the product of arithmetic averages of density and velocity: $\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,$. An investigation of other choices of $\mathcal{F}_{\rho}$, all leading to KEP EC schemes that could be better optimized for further targets, is postponed to future research.
@@ 4.2 Ideal gas
where $\overline{\phi}^{\text{log}}=\delta^{\,+}\phi/\delta^{\,+}\log\phi$ is the logarithmic mean, $\overline{\overline{pu}}\,=\left(p_{i}u_{i+1}+p_{i+1}u_{i}\right)$ is the product mean (corresponding with the FD approximation of $p\partial u/\partial x+u\partial p/\partial x$, representing the advective form of the pressure term in the total-energy equation), and the apex is used to indicate that the flux includes also the contribution coming from the (conservative) pressure term.
@@
^$\hat{\mathcal{F}}_{\rho e}=\dfrac{\gamma P}{\gamma-1}\hat{\mathcal{F}}_{\rho}\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}\rho}-UP=\text{const.}\quad\implies\quad\hat{\mathcal{F}}_{\rho}\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}\rho}=\text{const($U$,$P$)},$ (23)
@@
^$\hat{\mathcal{F}}_{\rho}=\dfrac{\delta^{\,+}\rho}{\delta^{\,+}\log\rho}U=\overline{\rho}^{\text{log}}U.$ (24)
@@
$\mathcal{F}_{\rho}=\dfrac{\delta^{\,+}\rho u}{\delta^{\,+}\log\rho}$ (25)

### v2

@@ Abstract
This study proposes a novel spatial discretization procedure for the compressible Euler equations which guarantees entropy conservation at a discrete level when an arbitrary equation of state is assumed. The proposed method, based on a locally-conservative discretization, guarantees also the spatial conservation of mass, momentum, and total energy and is kinetic-energy-preserving. In order to achieve the entropy-conservation property for an arbitrary non-ideal gas, a general strategy is adopted based on the manipulation of discrete balance equations through the imposition of global entropy conservation and the use of a summation by parts rule. The procedure, which is extended to an arbitrary order of accuracy, conducts to a general form of the internal-energy numerical flux which results in a nonlinear function of thermodynamic and dynamic variables and still admits the mass flux as a residual degree of freedom. The effectiveness of the novel entropy-conservative formulation is demonstrated through numerical tests making use of some of the most popular cubic equations of state.
@@ 1 Introduction
Numerical simulations are key in understanding the physical characteristics of turbulent compressible flows and are an invaluable tool for application-oriented studies [1]. However, it is well known that, even in the absence of shock waves, standard discretizations are susceptible to strong nonlinear instabilities at high values of the Reynolds number, especially when non-dissipative schemes are employed for the discretization of the convective terms [2]. This behavior can be traced back to the failure of the numerical method to reproduce at a discrete level some structural properties or invariants of the continuous system of equations. In fact, the compressible Euler equations, which are a mathematical expression of the conservation of mass, momentum and total energy, have many induced symmetries which are typically not reproduced by the discrete system. The numerical community has progressively realized that the traditional focus on classical concepts of numerical analysis, such as order of accuracy and asymptotic convergence, has to be complemented by some more ‘physical’ concepts, such as the discrete reproduction of the basic structural properties of the continuous formulation on an arbitrary finite mesh. These methods are referred to as structure preserving or physics compatible methods and have been the subject of many studies in recent years [2, 3, 4, 5].
@@
Examples of structural properties for the compressible flow equations are the conservation of entropy (for smooth, nonviscous flows), the conservation of kinetic energy (for nonviscous flows in the incompressible limit), and the equilibrium of constant pressure and velocity distributions, which allows the existence of density waves moving at constant velocity and pressure. Several numerical methods have been developed, which are able to preserve these properties at discrete level. Kinetic Energy Preserving (KEP) methods are the most well-known and studied among the others. These methods are designed to automatically preserve the conservative structure of the convective terms in the induced discrete equation for the evolution of kinetic energy. The study of KEP methods for compressible flows was pioneered by Feiereisen et al. [6]. Inspired by some formulations used for incompressible flows, the authors proposed for the first time the use of a skew-symmetric-like splitting of the convective terms in mass and momentum equations to induce a conservative structure of the convective terms in the kinetic-energy equation. This work stimulated the attention on this important topic and several other studies have appeared over the years, among which we mention here the Finite Volume (FV) analyses of Jameson [7] and Subbareddy and Candler [8] and the works on the triple splitting by Kennedy and Gruber [9] and Pirozzoli [10]. More recently, also less conventional approaches have been explored, in which square-root variables are used [11, 12, 13, 14]. A quite complete characterization of the possible KEP formulations in a Finite-Difference (FD) framework has been derived only recently [15], and its relations with FV formulations have also been explored [16, 17]. As KEP methods involve a proper discretization of mass and momentum equations alone, they are agnostic to the gas model used in the integration of the total-energy equation. KEP methods have been applied mainly in the context of ideal gases, but applications to more general equations of state have shown improved robustness as well [18, 19].
@@
Discrete entropy conservation has also been explored to improve fidelity and suppress instabilities in turbulent simulations. Most of the contributions have been developed by following the elegant theory proposed by Tadmor [20, 21] in the context of FV methods. Entropy Conservative (EC) numerical fluxes have been derived since the early papers on this subject in the context of ideal gases [20, 21, 22]. Ismail and Roe [23] developed ‘affordable’ EC fluxes based on the logarithmic mean, whereas Chandrashekar [24] proposed for the first time both KEP and EC numerical fluxes. More recently, Ranocha [25, 26, 27] proposed a KEP and EC formulation which is also able to enforce the so-called Pressure Equilibrium Preserving (PEP) property ([28, 27, 18, 29]), i.e. the ability of the discrete scheme to reproduce the traveling density wave solutions of the compressible Euler equations obtained for initially uniform pressure and velocity distributions. Recently, De Michele and Coppola [30] proposed a family of Asymptotically Entropy Conservative (AEC) schemes which are PEP and KEP, and arbitrarily reduce the error on entropy conservation. The main advantage of these schemes is that, being based on arithmetic and harmonic means, they only use algebraic operations for the computation of the fluxes, and do not exhibit singularities, with increased efficiency at comparable performances. AEC schemes improve already existing formulations which approximately enforce the entropy-conservation property, as the KEEP or KEEP(N) schemes [31, 32].
An alternative approach to the theory of Tadmor to produce EC scheme is represented by the residual distribution schemes proposed by Abgrall [33]. This correction procedure is not limited to the EC property, but it allows the incorporation of additional conservation laws into a baseline method. For example, it has been used to enforce various secondary properties such as angular momentum conservation [34] and KEP [35]. Furthermore, thermodynamically compatible schemes—having the entropy inequality as the primary evolution equation—have been produced which are able to discretely conserve total energy [36], These approaches can be independent of the specific equation of state, however a comparison with the proposed methodology will not be attempted in this study and is left for future research.
@@
All the mentioned studies on formulations which are EC in the sense of Tadmor have been developed in the context of ideal gases, for which a simple linear relation between pressure and internal energy is assumed. However, numerical simulations of real gases, often in transcritical regimes, have grown significantly in recent years, due to a wide range of possible engineering applications [37, 38, 39, 19]. Despite this growing interest, the literature on structure-preserving methods explicitly designed for real gases is still scarce. Most of the studies have been focused on the pressure equilibrium property, as the generation of spurious pressure oscillations, especially in the case of multi-component flows, is a longstanding issue in scale-resolving simulations with real gases [40, 41, 42, 18]. Despite these efforts, no numerical method is known at present that can discretely preserve the primary invariants mass, momentum, and total energy and simultaneously enforce the pressure equilibrium property or the conservation of entropy in the case of gases with an arbitrary equation of state.
@@ 2 Problem formulation
where $\rho$ is the density, $p$ the pressure, $u_{\alpha}$ the Cartesian velocity component along the direction $x_{\alpha}$ and $E$ the total energy per unit mass, sum of internal and kinetic energies: $E=e+u_{\alpha}u_{\alpha}/2$. In this last equation, in Eqs. (1)–(3) and in what follows, the summation convention over repeated Greek indices is assumed. The system (1)–(3) is closed by the specification of an equation of state, which is assumed in the form $p=p(\rho,e)$ and is left as unspecified for the moment.
In this paper we focus on the discretization of the spatial terms of the system of equations (1)–(3) in the context of a semi-discretized procedure. As is usual in this framework, after the spatial discretization step has been performed, the resulting system of Ordinary Differential Equations (ODE) is integrated in time by using a standard solver, and we assume that the effects of time integration errors are negligible at sufficiently small time steps. Consequently, in our theoretical analysis we assume that all the manipulations involving time derivatives can be carried out at the continuous level. Note that, as in many other similar studies, we deal with secondary conservation properties of spatial discretizations based on central schemes, which are specifically designed for shock-free simulations and, due to the absence of dissipation, are suited to turbulent flows in typically under-resolved simulations. Of course, these methods can also be used in smooth regions of shocked flows, in the spirit of hybrid methods [43, 44], which rely on shock sensors restricting the use of dissipative shock-capturing methods to regions near the shocks.
Although we will repeatedly refer to numerical fluxes, which are usually specified in the context of FV methods, our framework is mainly that of a FD treatment of the spatial terms in Eqs. (1)–(3). Numerical fluxes arise in our analysis because we assume that the discrete formulation of the convective terms in Eqs. (1)–(3), which are directly discretized, is always locally conservative, i.e. it can be expressed as the sum of differences of numerical fluxes at adjacent faces along the three coordinate directions. A locally-conservative discretization is also globally conservative, i.e. the sum of all the pointwise contributions on the domain is zero (or a function of boundary conditions) by virtue of the telescoping property. These concepts are familiar in FV methods, but our numerical fluxes are always interpreted as functions of nodal values over a colocated uniform mesh, although a generalization of our findings to more general situations and to other discretization techniques can be easily carried out in most cases. Latin subscripts as $i,j$ or $k$ are used to denote the values of the discretized variable on a nodal point $x_{i}$.
@@
and subtraction of Eq. (4) from Eq. (3) gives the evolution equation for internal energy
@@ 3 Discrete formulation
where $\mathcal{C}_{\rho\phi}$ is the convective term and $\mathcal{P}_{\rho\phi}$ is a pressure term. In Eqs. (1)–(6) $\phi$ assumes the values $1,u_{\beta},E,\kappa,e$ and $s$, respectively, whereas the pressure term is absent in mass (1) and entropy (6) equations. The symbols $\mathcal{C}$ and $\mathcal{P}$ will be used indifferently to denote both the continuous or discrete versions of the spatial terms at the right-hand sides of Eqs. (1)–(6). When needed, we will use the more explicit notation $\left.\mathcal{C}_{\rho\phi}\right|_{i}$ to denote a particular discretization of $\mathcal{C}_{\rho\phi}$ at node $x_{i}$. The theory will be developed for the one-dimensional version of Eq. (1)–(3), the extension to the three-dimensional case being straightforward, since the convective terms along the various directions are independent and can be treated separately by components. The 1D formulation uses an ordered set of grid nodes, thus we use a structured Cartesian mesh for the extension to two and three dimensions, and for carrying out the numerical tests (see Sec. 7.1). Such theoretical setting is, in principle, applicable also to unstructured meshes, at least for second-order, two-point fluxes. However, a high-order formulation as the one presented in Sec. 6 should be redesigned due to the lack of ordering between the nodes if unstructured grids are used. In this section and in the following Sec. 4 we will expose the theory with respect to second-order formulations, corresponding to two-point fluxes. In Sec. 6 the extension to high-order fluxes for Cartesian meshes is carried on.
@@
where $h$ is the spatial step size and $\delta^{\,-}$ is the backward difference operator: $\delta^{\,-}a_{i}=a_{i}-a_{i-1}$. In what follows we will need to consider also the forward difference operator, which is denoted by $\delta^{\,+}$, with $\delta^{\,+}b_{i}=b_{i+1}-b_{i}$. The operators $\delta^{\,-}$ and $\delta^{\,+}$ are adjoints with respect to the standard inner product. This means that, neglecting the effects of boundary conditions (or by assuming periodic boundary conditions), the following property is easily seen to hold
@@ 4.1 General treatment
A KEP discretization of mass and momentum equations can be implemented with a latitude on the specification of the mass flux $\mathcal{F}_{\rho}$. Once this has been done, the kinetic-energy flux is fixed as in Eq. (10), and one is left with the internal-energy flux $\mathcal{F}_{\rho e}$ as additional residual degree of freedom to satisfy further structural properties of the formulation. In this section, we analyze the induced discrete equation for the specific entropy $\rho s$ to obtain a formulation that is globally conservative of entropy.
@@
where $g=e-Ts+p/\rho$ is the Gibbs free energy. This equation (also reported as Eq. (15.5.2) in [47], p. 352) reduces to $\partial_{t}(\rho s)=c_{v}/e\,\partial_{t}(\rho e)+(s-\gamma c_{v})\,\partial_{t}\,\rho$ (see [15, Eq. (33)]) in the case of ideal gases. Equation (12) is now used to obtain a relation between the discretized spatial terms of mass, internal energy and entropy. In fact, substituting the right-hand sides of Eq. (1), (5) and (6), as expressed in the generic form of Eq. (7), one has
@@
We now assume that mass and internal-energy convective terms are discretized using a locally-conservative formulation, as expressed in Eq. (8): $\mathcal{C}_{\rho}=\delta^{\,-}\mathcal{F}_{\rho}/h$ and $\mathcal{C}_{\rho e}=\delta^{\,-}\mathcal{F}_{\rho e}/h$. Moreover, for the pressure term in the internal-energy equation $\mathcal{P}_{\rho e}=p\partial u/\partial x$ we use a standard, second-order central derivative formula, which can be expressed as
@@
$\dfrac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho s\,\mathrm{d}x=-\int_{\Omega}\mathcal{C}_{\rho s}\,\mathrm{d}x=\text{b.t.}$ (16)
@@
$\sum_{i}\left.\mathcal{C}_{\rho s}\right|_{i}\,h=\sum_{i}\left(\dfrac{1}{T}\,\delta^{\,-}\mathcal{F}_{\rho e}-\dfrac{g}{T}\,\delta^{\,-}\mathcal{F}_{\rho}+\dfrac{p}{T}\,\delta^{\,-}\overline{u}\,\right)=\text{b.t.}$ (17)
@@
Eq. (20) is the fundamental relation defining the class of EC schemes we analyze in this paper which still contains the mass flux $\mathcal{F}_{\rho}$ as a residual degree of freedom, meaning that an arbitrary choice of the analytical flux discretization would still yield a valid EC scheme. In the present study we will only consider the scheme having as a mass flux simply the product of arithmetic averages of density and velocity: $\mathcal{F}_{\rho}=\overline{\rho}\,\overline{u}\,$. An investigation of other choices of $\mathcal{F}_{\rho}$, all leading to KEP EC schemes that could be better optimized for further targets, is postponed to future research.
@@ 4.2 Ideal gas
where $\overline{\phi}^{\text{log}}=\delta^{\,+}\phi/\delta^{\,+}\log\phi$ is the logarithmic mean, $\overline{\overline{pu}}\,=\left(p_{i}u_{i+1}+p_{i+1}u_{i}\right)/2$ is the product mean (corresponding to the FD approximation of $p\partial u/\partial x+u\partial p/\partial x$, representing the advective form of the pressure term in the total-energy equation), and the apex is used to indicate that the flux includes also the contribution coming from the (conservative) pressure term.
@@
$\hat{\mathcal{F}}_{\rho e}=\dfrac{\gamma P}{\gamma-1}\hat{\mathcal{F}}_{\rho}\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}\rho}-UP=\text{const.}\quad\implies\quad\hat{\mathcal{F}}_{\rho}\dfrac{\delta^{\,+}\log\rho}{\delta^{\,+}\rho}=\text{const($U$,$P$)},$ (23)
@@
$\hat{\mathcal{F}}_{\rho}=\dfrac{\delta^{\,+}\rho}{\delta^{\,+}\log\rho}U=\overline{\rho}^{\text{log}}U.$ (24)
@@
$\mathcal{F}_{\rho}=\dfrac{\delta^{\,+}\rho u}{\delta^{\,+}\log\rho u}.$ (25)